Jim can fill a pool carrying bucks of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?
Answer by fractalier(6550)

Here's the setup...
t/30 + t/45 + t/90 = 1
now multiply everything by 90 to clear fractions and then solve for t...
90[t/30 + t/45 + t/90 = 1]
3t + 2t + t = 90
6t = 90
t = 15 minutes